# biblio.bib Metaheuristics and Large Language Models Join Forces: Towards an Integrated Optimization Approach

### v1

@@ Abstract
Since the rise of Large Language Models (LLMs) a couple of years ago, researchers in metaheuristics (MHs) have wondered how to use their power in a beneficial way within their algorithms. This paper introduces a novel approach that leverages LLMs as pattern recognition tools to improve MHs. The resulting hybrid method, tested in the context of a social network-based combinatorial optimization problem, outperforms existing state-of-the-art approaches that combine machine learning with MHs regarding the obtained solution quality. By carefully designing prompts, we demonstrate that the output obtained from LLMs can be used as problem knowledge, leading to improved results. Lastly, we acknowledge LLMs’ potential drawbacks and limitations and consider it essential to examine them to advance this type of research further.
1 Introduction
@@
The advent of Large Language Models (LLMs) has altered the Natural Language Processing (NLP) landscape, empowering professionals across diverse disciplines with their remarkable ability to generate human-like text. Models like OpenAI’s GPT [openai2023gpt4], Meta’s Llama [touvron2023llama], and Anthropic’s Claude 3 [Anthropic2024-ee] have become indispensable collaborators in many peoples’ daily lives; giving rise to innovative products such as ChatGPT for general use, GitHub Copilot for code generation, DALL-E 2 for image creation, and a multitude of voice generators, including OpenAI’s text-to-speech API and ElevenLabs’s Generative Voice AI. Currently, LLMs are being experimentally applied across various fields, yielding mixed results [ALMEIDA2024104145]. While some applications seem questionable, others exhibit spectacular outcomes. One of the most contentious applications is using LLMs for tasks necessitating mathematical reasoning. Given LLMs’ inherently probabilistic nature, this application was once deemed implausible. However, recent findings suggest a shift in perspective, particularly with LLMs boasting vast parameter counts [ahn2024large]. As LLMs continue to scale, new capabilities emerge [wei2022emergent]. Crucially, these opportunities are contingent upon the thoughtful design of prompts, which helps mitigate the risk of LLMs providing irrelevant or inaccurate responses [wan2024efficient].
@@
Whenever a new technology emerges, it is natural to wonder if it might enhance an existing one. In combinatorial optimization, metaheuristics (MHs) [10.5555/1941310] have been established as effective approximate algorithms for tackling complex, NP-hard problems. While they excel in rapidly providing good-enough solutions, they depend heavily on domain-specific knowledge. To address this limitation, researchers have explored the integration of MHs with other approaches, including exact algorithms and Machine Learning (ML). While combining MHs with exact algorithms has shown promise [BLUM20114135], a successful integration demands significant technical expertise. Alternatively, incorporating ML techniques within MHs can provide valuable problem insights [KARIMIMAMAGHAN2022393], but these approaches often require specialized know-how or, in the case of Deep Learning (DL), substantial datasets and computational resources for training [barocas2023fairness]. This paper seeks another direction. We delve into the potential of LLMs to create a novel hybrid approach that combines the strengths of MHs and LLMs.
1.1 Our contribution
We introduce a novel approach to enhance the performance of MHs by using LLMs’ output. While existing research has already started to explore the integration of LLMs within MHs, our strategy takes a distinct pattern recognition approach. Unlike previous methods that utilize LLMs to develop new MHs (code generation, e.g., [10.1145/3583133.3596401, ma2024llamoco]) or solve optimization problems through natural language descriptions [yang2023large], our approach focuses on designing prompts that include problem instance metrics, graph metrics in our application example. Thus, we employ LLMs not as an oracle providing final answers (i.e., it is not an end-to-end approach according to the classification by Bengio et al. [bengio2020machine]) but as an intermediate step, assisting in pattern detection within the metrics’ values (see Figure 2), i.e., it is an hybrid one.
@@ 1.1 Our contribution
The paper is organized as follows. Section 2 examines existing approaches for integrating ML into MHs and provides an overview of existing research on applying LLMs in optimization. Section 3 formally defines the NP-Hard combinatorial optimization problem that serves as an example for our study. Our proposed integration strategy for combining MH and LLMs is presented in Section 4. The empirical evaluation of our hybrid approach is detailed in Section 5, where we introduce a comprehensive three-dimensional framework for assessment and provide a visual analysis of the algorithm’s performance. Section 6 identifies remaining open research questions and discusses the current limitations of LLMs. Finally, Section 7 concludes the paper by summarizing our key findings. Moreover, future research directions are mentioned.
2 Background
2.1 Machine Learning for Enhancing Metaheuristics in Combinatorial Optimization
@@
Metaheuristics (MHs) are approximate algorithms that have proven effective in solving complex optimization problems, especially combinatorial optimization problems (COPs). COPs are characterized by discrete variables and a finite search space. Although MHs have been shown to deliver good results in reduced computation times, they do not guarantee finding the optimal solution. Moreover, their success often hinges on the availability of problem-specific knowledge. Each problem instance is treated similarly, applying problem-specific heuristics and (generally) relying on a stochastic behavior. Along these lines, the community aims to innovate by integrating techniques from various domains to improve MHs’ performance and address the limitations of MH techniques. In particular, hybrid approaches based on the combination with (1) exact algorithms [BLUM20114135] and (2) learning techniques [KARIMIMAMAGHAN2022393] have been explored. Currently, the primary focus has shifted towards the second option. Especially the integration of machine learning (ML) techniques has recently resulted in a multitude of different hybrid approaches. Researchers have explored various strategies to integrate ML into MHs [KARIMIMAMAGHAN2022393]. In particular, ML might be used for the following purposes in MHs: algorithm selection (determining the best MH for a given problem), fast approximate fitness evaluation in the presence of costly objective functions, initialization (generating high-quality initial solutions), and parameter configuration (optimizing the numerous parameters of an MH, which is crucial for its performance). These strategies utilize learning techniques to analyze numerous cases and scenarios, uncovering hidden patterns in the data. By identifying these patterns, MHs can extract general principles that apply to a broad range of situations. This enhances MH’s decision-making process, increasing their performance and adaptability.111The reverse integration, which involves enhancing ML architectures with MH techniques, is beyond the scope of this study.
@@
MH+ML: In a study by Sun et al. [SUN2022105769], ML techniques were incorporated into the metaheuristic Ant Colony Optimization (ACO) to address the orienteering problem. The authors improved the solution construction process of ACO by utilizing guided predictions based on engineered features. In [10.1007/978-3-030-53552-0_15], the authors developed a novel algorithm that combines a metaheuristic with decision trees to address the classic vehicle routing problem.
@@
MH+DL: Examples of this type of hybridization can be found abundantly in the literature of recent years. For instance, in [9909110], the authors presented a novel approach that uses a Graph Neural Network (GNN) for learning heuristic information that is then used by a Biased Random Key Genetic Algorithm (BRKGA) to translate random keys into solutions to the tackled problem. Another example concerns [liu2023machine] where the authors apply different GNN architectures for parametrizing the neighbor selection strategy in Tabu Search (TS) and in Large Neighborhood Search (LNS).
@@
MH+RL: RL has recently been used in numerous hybrid approaches. For instance, in [huber2021learning], the authors devised a method for learning the heuristic function of beam search in the context of two variants of the Longest Common Subsequence (LCS) problems. In [FENOY2024104064], the authors proposed a hybrid approach comprised of an attention-based model trained with RL and combined with a more classical optimization method for the formation of collectives of agents in real-world scenarios, showing that it reaches the performance of state-of-the-art solutions while being more general. Furthermore, a variable neighborhood search (VNS) based on Q-learning was devised for a machine scheduling problem in [alicastro2021reinforcement]. Finally, we also mention [chaves2021adaptive], where RL is used for adapting the parameters of a BRKGA during the evolutionary process.
@@
While these hybridization strategies offer potential solutions, they each come with their own set of drawbacks. For instance, the manual feature selection process in many MH+ML approaches relies heavily on the expertise of a specialist. Concerning MH+DL approaches, the system’s generalization ability might be hindered by lacking a large and diverse enough dataset. As for MH+RL hybrids, the complexity lies in defining the action space, rewards, and learning policies in a clear and effective manner. Moreover, all three approaches share similar technical challenges: the complexity of replicating models, the time-consuming process of data collection and preparation, and the computational demands of generalization, especially for large-scale problems or complex optimization tasks [barocas2023fairness].
@@
To explore innovative methods for enhancing the performance of MHs and considering the usefulness and potential of LLMs, which will be discussed in the following subsection, in this paper, we explore a novel hybrid approach: MH+LLM (see Section 4). Utilizing the capabilities of LLMs—while being aware of their limitations—we aim to enhance the problem-solving capabilities of MHs and open up new avenues for tackling complex optimization problems.
2.2 LLMs as Pattern Recognition Engines
@@
LLMs have breathed new life into the field of NLP. These high-level language models employ billions of parameters and exhibit an outstanding ability to learn from data. Models like GPT-4o (OpenAI)222https://openai.com/index/hello-gpt-4o., Claude-3-Opus (Anthropic)333https://www.anthropic.com/news/claude-3-family., Gemini 1.5 (Google)444https://deepmind.google/technologies/gemini., Mixtral 8x22b (Mistral AI)555https://mistral.ai/news/mixtral-8x22b., and Command-R+ (Cohere)666https://docs.cohere.com/docs/command-r-plus., as well as tools built on top of them—such as ChatGPT, GitHub Copilot, and Bing Chat—have demonstrated that we are in the presence of a groundbreaking technology. Unlike previous advancements, this new wave of AI is no longer limited to experts; instead, it is accessible to anyone who can grasp its benefits.
@@
LLMs are generative AI models that produce text sequentially, predicting each token based on the previous ones. This is made possible by the Transformer, a groundbreaking DL architecture that revolutionized the field of NLP. Proposed by Vaswani et al. [vaswani2023attention], it introduces the concept of self-attention, allowing the model to contextually select the most suitable words. The Transformer derives its name from its ability to transform a set of vectors in a given representation space to a new set of vectors with identical dimensions but in a different space. By assigning varying weight values to each input, the attention mechanism leverages inductive biases related to sequential data [M2023-wp]. This architecture also takes advantage of the capabilities of high-performance hardware due to its parallelizable nature. As a result, the Transformer generates words that seamlessly fit the context—although it lacks factual verification—marking a significant advancement in NLP tasks.
@@
LLMs have found uses in various domains, including the interpretation of complex results like chemical compounds or images [singh2024rethinking], where they provide insights and explanations. Additionally, LLMs are being applied as autonomous agents that leverage external tools and resources to accomplish tasks [guo2024large]. Furthermore, these models demonstrate progress in domains once thought to be out of the reach of their capabilities, including mathematical reasoning and optimization tasks [yang2023large]. While there are still numerous obstacles to overcome [ahn2024large], these advancements showcase the potential of these models to address intricate cognitive problems.
@@
Recent research on leveraging LLMs for optimization has primarily explored two paths: first, formulating optimization problems within the prompt and requesting the LLM to solve the described problem, typically for straightforward optimization tasks [yang2023large]; and second, automating code generation to enhance optimization algorithms [ma2024llamoco, 10.1145/3583133.3596401]. Nevertheless, a method that harnesses LLMs as pattern recognition engines in combinatorial optimization problems has not yet been developed, even though a recent study has demonstrated the ability of LLMs to find patterns in various tasks [mirchandani2023large]. Our proposal introduces an integration that enhances the effectiveness of the metaheuristic search process by utilizing LLMs, marking a step forward in that direction. While acknowledging the challenges associated with LLMs (discussed in the subsequent Section 2.2.1), we identify numerous opportunities for progress in this area (see Section 4).
2.2.1 Obstacles and Opportunities
@@
LLMs have no built-in understanding of the world, as they cannot directly experience or “simulate” our environment. This is unlike humans, who constantly perceive information, be it sensory, visual, or auditory. In contrast, LLM operation depends on the data with which they are trained. Consequently, if the data is not meticulously selected by humans, the model’s predictions may be unreliable or even generate false information, a phenomenon known as “hallucinations” in AI discourse.777The term “hallucination” in the context of LLMs was derived from the concept of “AI hallucination,” which refers to incorrect responses generated by AI systems. In fact, this term was adopted due to the tendency of humans to anthropomorphize technology, attributing human qualities to it. As Floridi and Nobre have recently shown [Floridi2024-jv], such a tendency is common with disruptive technologies. It has also been shown that, as we become more familiar with these technologies, the tendency to anthropomorphize should decrease over time. We mitigated the problem of hallucinations with a special focus on prompt design. It has been observed that ambiguous prompts can lead to inconsistent results, with the same prompt potentially yielding different responses each time it is executed. As LLMs have evolved and increased in size, it has become evident that altering the prompt strategy can significantly improve the quality of the responses. This discovery has opened up new possibilities, enabling LLMs to tackle tasks that previously yielded negative or untested results. Carefully designing and adjusting prompts can improve the reliability and performance of LLMs [qiao2023reasoning]. This article focuses on creating prompts that produce desired outcomes.
3 Problem Definition
In this section, we provide the definition of the optimization problem we consider as an example in this paper, while in Section 4, we present our hybrid optimization approach. The considered optimization problem is from the realm of social networks. In fact, social network problems often serve as an interesting experimental laboratory for testing optimization techniques since they can be modeled as combinatorial problems based on directed graphs that become highly complex as the instance size grows.
Specifically, we consider Multi-Hop Influence Maximization, a social network problem proven NP-hard by Ni et al. [Ni] and Basuchowdhuri et al. [Basuchowdhuri]. In the literature, this problem has been studied using both metaheuristics and a state-of-the-art combination of a BRKGA with deep learning (DL) [9909110], which gives us a point of comparison.
3.1 Multi-Hop Influence Maximization in Social Networks

### v2

# Metaheuristics and Large Language Models Join Forces: Towards an Integrated Optimization Approach

@@ Abstract
Since the rise of Large Language Models (LLMs) a couple of years ago, researchers in metaheuristics (MHs) have wondered how to use their power in a beneficial way within their algorithms. This paper introduces a novel approach that leverages LLMs as pattern recognition tools to improve MHs. The resulting hybrid method, tested in the context of a social network-based combinatorial optimization problem, outperforms existing state-of-the-art approaches that combine machine learning with MHs regarding the obtained solution quality. By carefully designing prompts, we demonstrate that the output obtained from LLMs can be used as problem knowledge, leading to improved results. Lastly, we acknowledge LLMs’ potential drawbacks and limitations and consider it essential to examine them to advance this type of research further. Our method can be reproduced using a tool available at: https://github.com/camilochs/optipattern.
Index Terms:
combinatorial optimization, hybrid algorithm, metaheuristics, large language models=-15pt
I Introduction
@@
The advent of Large Language Models (LLMs) has altered the Natural Language Processing (NLP) landscape, empowering professionals across diverse disciplines with their remarkable ability to generate human-like text. Models like OpenAI’s GPT [1], Meta’s Llama [2], and Anthropic’s Claude 3 [3] have become indispensable collaborators in many peoples’ daily lives; giving rise to innovative products such as ChatGPT for general use, GitHub Copilot for code generation, DALL-E 2 for image creation, and a multitude of voice generators, including OpenAI’s text-to-speech API and ElevenLabs’s Generative Voice AI. Currently, LLMs are being experimentally applied across various fields, yielding mixed results [4]. While some applications seem questionable, others exhibit spectacular outcomes. One of the most contentious applications is using LLMs for tasks necessitating mathematical reasoning. Given LLMs’ inherently probabilistic nature, this application was once deemed implausible. However, recent findings suggest a shift in perspective, particularly with LLMs boasting vast parameter counts [5]. As LLMs continue to scale, new capabilities emerge [6]. Crucially, these opportunities are contingent upon the thoughtful design of prompts, which helps mitigate the risk of LLMs providing irrelevant or inaccurate responses [7].
@@
Whenever a new technology emerges, it is natural to wonder if it might enhance an existing one. In combinatorial optimization, metaheuristics (MHs) [8] have been established as effective approximate algorithms for tackling complex, NP-hard problems. While they excel in rapidly providing good-enough solutions, they depend heavily on domain-specific knowledge. To address this limitation, researchers have explored the integration of MHs with other approaches, including exact algorithms and Machine Learning (ML). While combining MHs with exact algorithms has shown promise [9], a successful integration demands significant technical expertise. Alternatively, incorporating ML techniques within MHs can provide valuable problem insights [10], but these approaches often require specialized know-how or, in the case of Deep Learning (DL), substantial datasets and computational resources for training [11]. This paper seeks another direction. We delve into the potential of LLMs to create a novel hybrid approach that combines the strengths of MHs and LLMs.
I-A Our contribution
We present a novel approach to enhance MHs performance utilizing LLMs’ output. Our method diverges from existing techniques in two key aspects:
Unlike methods that use LLMs as direct optimizers for natural language-described problems [15, 16, 17]—a technique limited to simple optimization tasks due to LLMs’ stochastic nature—our approach tackles complex optimization challenges by using LLMs to identify and track pertinent information within the problem instance.
This dual-faceted approach represents a significant advancement in the integration of LLMs with metaheuristic optimization techniques, offering a more robust and versatile framework for tackling a wide array of optimization challenges. Thus, we employ LLMs not as an oracle providing final answers (i.e., it is not an end-to-end approach according to the classification by Bengio et al. [18]) but as an intermediate step, assisting in pattern detection within the metrics’ values (see Figure IV), i.e., it is an hybrid one.
@@ I-A Our contribution
The paper is organized as follows. Section II examines existing approaches for integrating ML into MHs and provides an overview of existing research on applying LLMs in optimization. Section III formally defines the NP-Hard combinatorial optimization problem that serves as an example for our study. Our proposed integration strategy for combining MH and LLMs is presented in Section IV. The empirical evaluation of our hybrid approach is detailed in Section V, where we introduce a comprehensive three-dimensional framework for assessment and provide a visual analysis of the algorithm’s performance. Section VI identifies remaining open research questions and discusses the current limitations of LLMs. Finally, Section VII concludes the paper by summarizing our key findings. Moreover, future research directions are mentioned.
I-B Reproducibility
Recognizing the importance of reproducibility in our field [20], and the potential challenges introduced by new technologies, we have developed a tool called OptiPattern (LLM-Powered Pattern Recognition for Combinatorial Optimization) that automates our hybridization method—detailed in Section IV—to ensure greater ease and accuracy in replication. This tool allows researchers to input a problem instance, generating the full prompt in response. Furthermore, by incorporating an LLM API key, the tool can return node-specific probabilities, which can then be integrated into the metaheuristic.111https://github.com/camilochs/optipattern This is essential, as replicating prompts can be complex and prone to errors.
II Background
II-A Machine Learning for Enhancing Metaheuristics in Combinatorial Optimization
@@
Metaheuristics (MHs) are approximate algorithms that have proven effective in solving complex optimization problems, especially combinatorial optimization problems (COPs). COPs are characterized by discrete variables and a finite search space. Although MHs have been shown to deliver good results in reduced computation times, they do not guarantee finding the optimal solution. Moreover, their success often hinges on the availability of problem-specific knowledge. Each problem instance is treated similarly, applying problem-specific heuristics and (generally) relying on a stochastic behavior. Along these lines, the community aims to innovate by integrating techniques from various domains to improve MHs’ performance and address the limitations of MH techniques. In particular, hybrid approaches based on the combination with (1) exact algorithms [9] and (2) learning techniques [10] have been explored. Currently, the primary focus has shifted towards the second option. Especially the integration of machine learning (ML) techniques has recently resulted in a multitude of different hybrid approaches. Researchers have explored various strategies to integrate ML into MHs [10]. In particular, ML might be used for the following purposes in MHs: algorithm selection (determining the best MH for a given problem), fast approximate fitness evaluation in the presence of costly objective functions, initialization (generating high-quality initial solutions), and parameter configuration (optimizing the numerous parameters of an MH, which is crucial for its performance). These strategies utilize learning techniques to analyze numerous cases and scenarios, uncovering hidden patterns in the data. By identifying these patterns, MHs can extract general principles that apply to a broad range of situations. This enhances MH’s decision-making process, increasing their performance and adaptability.222The reverse integration, which involves enhancing ML architectures with MH techniques, is beyond the scope of this study.
@@
MH+ML: In a study by Sun et al. [21], ML techniques were incorporated into the metaheuristic Ant Colony Optimization (ACO) to address the orienteering problem. The authors improved the solution construction process of ACO by utilizing guided predictions based on engineered features. In [22], the authors developed a novel algorithm that combines a metaheuristic with decision trees to address the classic vehicle routing problem.
@@
MH+DL: Examples of this type of hybridization can be found abundantly in the literature of recent years. For instance, in [19], the authors presented a novel approach that uses a Graph Neural Network (GNN) for learning heuristic information that is then used by a Biased Random Key Genetic Algorithm (BRKGA) to translate random keys into solutions to the tackled problem. Another example concerns [23] where the authors apply different GNN architectures for parametrizing the neighbor selection strategy in Tabu Search (TS) and in Large Neighborhood Search (LNS).
@@
MH+RL: RL has recently been used in numerous hybrid approaches. For instance, in [24], the authors devised a method for learning the heuristic function of beam search in the context of two variants of the Longest Common Subsequence (LCS) problems. In [25], the authors proposed a hybrid approach comprised of an attention-based model trained with RL and combined with a more classical optimization method for the formation of collectives of agents in real-world scenarios, showing that it reaches the performance of state-of-the-art solutions while being more general. Furthermore, a variable neighborhood search (VNS) based on Q-learning was devised for a machine scheduling problem in [26]. Finally, we also mention [27], where RL is used for adapting the parameters of a BRKGA during the evolutionary process.
@@
While these hybridization strategies offer potential solutions, they each come with their own set of drawbacks. For instance, the manual feature selection process in many MH+ML approaches relies heavily on the expertise of a specialist. Concerning MH+DL approaches, the system’s generalization ability might be hindered by lacking a large and diverse enough dataset. As for MH+RL hybrids, the complexity lies in defining the action space, rewards, and learning policies in a clear and effective manner. Moreover, all three approaches share similar technical challenges: the complexity of replicating models, the time-consuming process of data collection and preparation, and the computational demands of generalization, especially for large-scale problems or complex optimization tasks [11].
@@
To explore innovative methods for enhancing the performance of MHs and considering the usefulness and potential of LLMs, which will be discussed in the following subsection, in this paper, we explore a novel hybrid approach: MH+LLM (see Section IV). Utilizing the capabilities of LLMs—while being aware of their limitations—we aim to enhance the problem-solving capabilities of MHs and open up new avenues for tackling complex optimization problems.
II-B LLMs as Pattern Recognition Engines
@@
LLMs have breathed new life into the field of NLP. These high-level language models employ billions of parameters and exhibit an outstanding ability to learn from data. Models like GPT-4o (OpenAI)333https://openai.com/index/hello-gpt-4o., Claude-3-Opus (Anthropic)444https://www.anthropic.com/news/claude-3-family., Gemini 1.5 (Google)555https://deepmind.google/technologies/gemini., Mixtral 8x22b (Mistral AI)666https://mistral.ai/news/mixtral-8x22b., and Command-R+ (Cohere)777https://docs.cohere.com/docs/command-r-plus., as well as tools built on top of them—such as ChatGPT, GitHub Copilot, and Bing Chat—have demonstrated that we are in the presence of a groundbreaking technology. Unlike previous advancements, this new wave of AI is no longer limited to experts; instead, it is accessible to anyone who can grasp its benefits.
@@
LLMs are generative AI models that produce text sequentially, predicting each token based on the previous ones. This is made possible by the Transformer, a groundbreaking DL architecture that revolutionized the field of NLP. Proposed by Vaswani et al. [28], it introduces the concept of self-attention, allowing the model to contextually select the most suitable words. The Transformer derives its name from its ability to transform a set of vectors in a given representation space to a new set of vectors with identical dimensions but in a different space. By assigning varying weight values to each input, the attention mechanism leverages inductive biases related to sequential data [29]. This architecture also takes advantage of the capabilities of high-performance hardware due to its parallelizable nature. As a result, the Transformer generates words that seamlessly fit the context—although it lacks factual verification—marking a significant advancement in NLP tasks.
@@
LLMs have found uses in various domains, including the interpretation of complex results like chemical compounds or images [30], where they provide insights and explanations. Additionally, LLMs are being applied as autonomous agents that leverage external tools and resources to accomplish tasks [31]. Furthermore, these models demonstrate progress in domains once thought to be out of the reach of their capabilities, including mathematical reasoning and optimization tasks [15]. While there are still numerous obstacles to overcome [5], these advancements showcase the potential of these models to address intricate cognitive problems.
@@
Recent research on leveraging LLMs for optimization has primarily explored two paths: first, formulating optimization problems within the prompt and requesting the LLM to solve the described problem, typically for straightforward optimization tasks [15, 16, 17]; and second, automating code generation to enhance optimization algorithms [12, 13, 14]. While both strategies are effective in leveraging LLMs to address certain weaknesses of MHs, they fail to account for the significance of problem instances, as different instances can produce varying results in an MH. Our approach seeks to tackle this challenge, functioning as a complementary tool rather than a rival to current strategies. Thus, we present a novel integration that utilizes LLMs to enhance the effectiveness of metaheuristic search processes. Although a recent study has shown that LLMs can detect patterns across various tasks [32], no method has yet been developed that employs LLMs as pattern recognition engines in combinatorial optimization problems. Acknowledging the challenges related to LLMs (discussed in the subsequent Section II-B1), we identify numerous opportunities for progress in this area (see Section IV).
II-B1 Obstacles and Opportunities
@@
LLMs have no built-in understanding of the world, as they cannot directly experience or “simulate” our environment. This is unlike humans, who constantly perceive information, be it sensory, visual, or auditory. In contrast, LLM operation depends on the data with which they are trained. Consequently, if the data is not meticulously selected by humans, the model’s predictions may be unreliable or even generate false information, a phenomenon known as “hallucinations” in AI discourse.888The term “hallucination” in the context of LLMs was derived from the concept of “AI hallucination,” which refers to incorrect responses generated by AI systems. In fact, this term was adopted due to the tendency of humans to anthropomorphize technology, attributing human qualities to it. As Floridi and Nobre have recently shown [33], such a tendency is common with disruptive technologies. It has also been shown that, as we become more familiar with these technologies, the tendency to anthropomorphize should decrease over time. We mitigated the problem of hallucinations with a special focus on prompt design. It has been observed that ambiguous prompts can lead to inconsistent results, with the same prompt potentially yielding different responses each time it is executed. As LLMs have evolved and increased in size, it has become evident that altering the prompt strategy can significantly improve the quality of the responses. This discovery has opened up new possibilities, enabling LLMs to tackle tasks that previously yielded negative or untested results. Carefully designing and adjusting prompts can improve the reliability and performance of LLMs [34]. This article focuses on creating prompts that produce desired outcomes.
III Problem Definition
In this section, we provide the definition of the optimization problem we consider as an example in this paper, while in Section IV, we present our hybrid optimization approach. The considered optimization problem is from the realm of social networks. In fact, social network problems often serve as an interesting experimental laboratory for testing optimization techniques since they can be modeled as combinatorial problems based on directed graphs that become highly complex as the instance size grows.
Specifically, we consider Multi-Hop Influence Maximization, a social network problem proven NP-hard by Ni et al. [35] and Basuchowdhuri et al. [36]. In the literature, this problem has been studied using both metaheuristics and a state-of-the-art combination of a BRKGA with deep learning (DL) [19], which gives us a point of comparison.
[h](topskip=0pt, botskip=0pt, midskip=0pt)[width=0.9]figures/kdDSP.pdf Multi-hop influence process. The given directed graph consists of 11 nodes and 12 arcs, and the task is to solve the -DSP with . The example solution consists of two nodes: and (colored in green). The bottom row illustrates the concept of -hop coverage: when , nodes are 1-hop covered by ; when , nodes are 2-hop covered by ; and when , all remaining nodes in the graph are 3-hop covered by .
III-A Multi-Hop Influence Maximization in Social Networks